# Small Batch Size Training for Language Models: When Vanilla SGD Works, and Why Gradient Accumulation Is Wasteful

### v1

@@ Abstract
Conventional wisdom dictates that small batch sizes make language model pretraining and fine-tuning unstable, motivating gradient accumulation, which trades off the number of optimizer steps for a proportional increase in batch size. While it is common to decrease the learning rate for smaller batch sizes, other hyperparameters are often held fixed. In this work, we revisit small batch sizes all the way down to batch size one, and we propose a rule for scaling Adam hyperparameters to small batch sizes. We find that small batch sizes (1) train stably, (2) are consistently more robust to hyperparameter choices, (3) achieve equal or better per-FLOP performance than larger batch sizes, and (4) notably enable stable language model training with vanilla SGD, even without momentum, despite storing no optimizer state. Building on these results, we provide practical recommendations for selecting a batch size and setting optimizer hyperparameters. We further recommend against gradient accumulation unless training on multiple devices with multiple model replicas, bottlenecked by inter-device bandwidth.
@@
Figure 1: Small batch sizes are more robust to optimizer design and hyperparameter values. (a) Loss achieved by a transformer decoder-only language model with 30M active parameters trained on 600M tokens of FineWeb-Edu data using SGD, Adafactor, Adam, and Muon. At small batch sizes, all optimizers achieve similar loss; at large batch sizes, the gap between optimizers grows. (b) Loss achieved by GPT-2 (124M) trained on 2.5B tokens of FineWeb using AdamW when tuning the learning rate and $\beta_{1}$ hyperparameters for batch sizes 1 and 512. While both batch sizes achieve similar lowest loss, the smaller batch size is much more robust to the hyperparameter values, reducing the need for hyperparameter tuning. For batch size 512, we use the default hyperparameters from Brown et al. [1] and the nanoGPT public repository [^2] https://github.com/karpathy/nanoGPT ($\beta_{2}=0.95,w=0.1$). For batch size 1, we turn off weight decay and rescale $\beta_{2}$ to preserve the token half-life, resulting in $\beta_{2}=0.9999$.
@@ 1 Introduction
Large batch sizes are widely believed to improve the stability of language model training [2, 3, 4, 5]. As a consequence, sophisticated optimizers that perform well in large-batch training are increasingly standard practice [6, 7, 8, 9, 10, 11]. In fact, it is common to simulate batch sizes even larger than the maximum batch size that fits into device memory through gradient accumulation [5, 4, 12, 13].
@@
In small batch size pretraining experiments, one may observe loss spikes and heavy instability [14, 15, 16, 17]. While it is common to decrease the learning rate for smaller batch sizes, other hyperparameters, such as the decay rates for the first and second moments in Adam ($\beta_{1}$ and $\beta_{2}$), are often held fixed across batch sizes. We show that if instead of holding $\beta_{2}$ fixed, we hold the half-life of $\beta_{2}$ fixed (measured in number of tokens), stable training is possible all the way down to batch size one, as illustrated in Figure 1(a).
@@
When scaling hyperparameters in this way, we find that small batch sizes can confer computational advantages and greater robustness. In the small batch regime, we observe that the speed of convergence is less sensitive to optimizer hyperparameters like the learning rate, and momentum becomes less necessary, as we find in Figure 1(b). In contrast, large batch sizes require large learning rates to compensate for taking fewer steps at a given compute budget. Large step sizes in turn require that the optimizer makes predictions about the loss surface far away from the current iterate. We find that predicting such far-away parameter updates requires a sophisticated and carefully tuned optimizer. For very small batch sizes, we find that even vanilla stochastic gradient descent (SGD), with no momentum or weight decay, becomes competitive for training language models, in contrast to findings from recent work [18]. The ability to perform stable training without momentum can confer significant memory advantages, since we do not need to store the corresponding optimizer state, which can have a significant memory footprint. Based on our observations, we also revisit Adafactor, a memory-efficient variant of Adam, and show that it can be a compelling alternative to Adam in the small batch regime, enabling training of larger models in a memory-constrained setting.
@@
These findings could have a significant bearing on practice. Rather than perform gradient accumulation, or use the largest batch size that fits into device memory, it could be preferable to use smaller batch sizes that enable simpler procedures, computational and memory advantages, and require less tuning. Our findings suggest a best practice of using the smallest batch size that maximizes model throughput. We find that these prescriptions hold in both LLM pretraining and fine-tuning.
@@
It is common in convex optimization to use higher batch sizes as the loss converges to a minimum [2, 19]. Intuitively, the higher the gradient to gradient noise magnitude ratio, the smaller a batch size we should take. For example, if the gradient noise magnitude is close to zero, then we obtain roughly the same gradient estimate with a small batch size as a large batch size but pay far fewer floating-point operations (FLOPs), so small batch sizes are efficient. Smaller batch sizes allow us to take more steps for a fixed FLOP budget. As we converge to a minimum, that same ratio instead often tends to zero since the gradient norm vanishes while the gradient noise may not, in which case we may want a larger batch size. We hypothesize that language model training lives in the far-from-convergence regime where small batch sizes are efficient. Following compute-optimal scaling laws, we would not train language models into the convergence regime since we could achieve better performance by instead training a larger model on less data [20, 21].
@@
The paper is structured as follows: in Sections 2 and 3, we cover the background and related work for language model training. In Section 4, we demonstrate that small batch sizes not only perform on par with larger batch sizes for several optimizers that we carefully tune, but they also exhibit more robustness to optimizer and hyperparameter choices. We also show that the gap between vanilla SGD and more sophisticated optimizers shrinks in the small batch regime. Moreover, we derive scaling heuristics for Adam’s hyperparameters and demonstrate that they yield improved performance when transferred to new, larger-scale settings. Motivated by our findings, we provide practical prescriptions for training language models in Section 5. Namely, we recommend using the smallest batch size that maximizes model throughput and advise against gradient accumulation for most practitioners. Additionally, we show that Adafactor [22] achieves the best trade-off between memory use and performance when the batch size is small. We close with a discussion of open questions in Section 6.
@@ 2 Background: Optimization for Language Models
Stochastic gradient descent (SGD) [23, 24]. SGD updates model parameters by computing gradients on mini-batches of data as follows: $\theta_{t+1}=\theta_{t}-\eta g_{t}$, where $\theta_{t}$ are the model parameters at step $t$, $\eta$ is the learning rate, and $g_{t}=\nabla_{\theta}\mathcal{L}(\theta_{t})$ is the gradient of the loss function with respect to $\theta_{t}$ averaged over $B$ samples. We refer to $B$ as the batch size and consider the extreme case of $B=1$ in our experiments.
We show that although SGD is extremely simple, it can perform competitively when correctly tuned at small batch sizes. Momentum is often added to accumulate a moving average of past gradients, smoothing updates and accelerating convergence. However, we focus on vanilla SGD in our experiments to make the point that, with a small batch size, even momentum is unnecessary.
Adam [25]. Adam is an adaptive optimizer that maintains an exponential moving average (EMA) of the gradient and squared gradient, referred to as the first and second moments, and denoted as $m_{t}$ and $v_{t}$. The timescales of these moving averages are controlled by decay rates $\beta_{1}$ and $\beta_{2}$:
@@
Higher values of $\beta_{1}$ and $\beta_{2}$ place more weight on past gradients, meaning that the moments are averaged over longer timescales and evolve slowly. When training language models, values like $\beta_{1}=0.9$ and $\beta_{2}\in[0.95,0.98]$ are often used in practice [1, 26]. While most researchers and practitioners do not scale $\beta_{1}$ and $\beta_{2}$ with the batch size, we show in our work that scaling $\beta_{2}$ in particular is crucial for achieving good performance with Adam at a small batch size.
@@
Moment half-life. We found it helpful across many of our experiments to measure the typical timescales (measured in number of tokens) that the first and second moment in Adam are averaged over, instead of directly working with $\beta_{1}$ and $\beta_{2}$.
At each Adam update step [eq. 1], the contribution of previous gradients to the first moment gets reduced by a factor of $\beta_{1}$ and the contribution to the second moment gets reduced by a factor of $\beta_{2}$. Hence by definition there exists a certain number of steps $n$, after which the contribution of any mini-batch gradient is halved: $\beta^{n}=\frac{1}{2}$. Since every update step corresponds to observing $(B\cdot T)$ tokens, where $B$ is the batch size and $T$ is the sequence length, we can convert the number of optimizer steps into a number of observed tokens. As such, we can express the number of tokens it takes to decay a mini-batch gradient by a factor of $\frac{1}{2}$ as the half-life of the decay rate, denoted by $t_{\nicefrac{{1}}{{2}}}$, where $\beta^{\frac{t_{\nicefrac{{1}}{{2}}}}{BT}}=\frac{1}{2}$. The half-life provides a measure of the “typical” timescale that gradients are averaged over. For example, if we use a half-life of $t_{\nicefrac{{1}}{{2}}}=10\textrm{M}$ tokens, this means that the second moment is averaged over roughly 10M tokens. Throughout this paper, we will refer to the half-life of the first moment as $t_{1}$ and the half-life of the second moment as $t_{2}$, analogously to the more conventional parameterization $\beta_{1}$ and $\beta_{2}$.
Adafactor [22]. The second moment estimate $v_{t}$ used in Adam requires storing a number of floating point values equal to the parameter count of the model itself. Instead, Adafactor only stores the per-row and per-column sums of the moving average for the neural network weight matrices, and estimates the per-parameter second moment using these sums.
In contrast with Adam, this implementation of Adafactor not only factorizes the second moment estimate but also does not store a first moment estimate at all. Therefore, the memory requirements for the moving averages are reduced from $2\times d_{1}\times d_{2}$ for Adam to $d_{1}+d_{2}$ for Adafactor. We demonstrate in Sections 4 and 5 that thanks to this sublinear memory cost, Adafactor achieves the best performance–memory trade-off.
Muon [10]. Another optimizer that was recently demonstrated to be competitive for language model training is MomentUm Orthogonalized by Newton-Schulz (Muon). Muon is specifically designed for $\geq 2$-dimensional parameter tensors, such as weight matrices in linear layers, where the Newton-Schulz iterative method is applied to the first moment estimate $m_{t}$ before using it to update the parameters. Muon is often used in conjunction with other optimizers like Adam. In our experiments using Muon, we only apply it to hidden layers and use Adam for input embedding and final layers.
@@ 3 Related Work
Batch size has long been studied in both theory and practice: its impact on the optimization path [27, 28, 29, 30], its interaction with different optimizers [31, 32, 33, 34], and its effect on the optimal learning rates and momentum hyperparameters [35, 36, 37, 38]. Prior work has also explored related themes, such as the flatness argument favoring SGD [39, 40, 41, 42], SGD’s implicit biases [43, 44, 45, 46], and how stochastic versus full-batch optimization compare [47, 48, 49, 50]. However, modern language model training is neither convex nor characterized by multiple epoch training like vision settings. In fact, we often train for a limited token budget and stop far before convergence, since otherwise scaling laws suggest that we could train a much better language model for the same compute budget by training a larger model on fewer tokens [21]. Hence, prior intuitions, for instance, such as the need for larger batch size in the convergence phase, may not apply. For this reason, in the rest of this section we cover works that focus on modern language model training specifically.
@@
Zhang et al. [51] and Shallue et al. [33] examine the critical batch size, which represents the threshold beyond which greater data parallelism would lead to diminishing returns, and tune all optimizer hyperparameters. They argue that any batch size below the critical batch size should perform equally well when training for the same number of tokens. Differently from these works, we empirically demonstrate that small batch sizes, including the extreme value of batch size one, perform on par with or even better than larger batch sizes, and we also show that small batches are more robust to optimizer and hyperparameter choices and enable SGD to perform competitively. We validate these findings on language models with modern design choices and up to 1.3 billion parameters.
@@
On the other hand, Vyas et al. [52] argue that gradient noise from small batch sizes does not yield similar performance advantages in the single epoch setting as it does in multiple-epoch training. Experimental results from Filatov et al. [53] also suggest that small batches do not work as well as large batches for language models trained on C4 [54], but notably the authors do not tune the decay rates $\beta_{1}$ and $\beta_{2}$ of the Adam optimizer. In the same vein, Xiao [15] claim that both excessively small and large batch sizes underperform for language model training with Adam, leading to a U-shaped curve of the loss as a function of the batch size. We reproduce the exact experiments performed by Xiao [15] and show that their conclusions are an artifact of incomplete hyperparameter tuning, and tuning the decay rate of the second moment estimate alone makes small batch sizes competitive.
@@
Zhao et al. [18] train language models on C4 using different diagonal preconditioning optimizers, namely Adam, Lion, Adafactor with momentum, SGD with momentum, and signed SGD with momentum. They demonstrate that all these optimizers, aside from SGD which lags behind, achieve comparable optimal performance and robustness to hyperparameters. We challenge this conclusion in our work and argue that in the small batch size regime, SGD can perform on par with Adam.

### v2

@@ Abstract
Conventional wisdom dictates that small batch sizes make language model pretraining and fine-tuning unstable, motivating gradient accumulation, which trades off the number of optimizer steps for a proportional increase in batch size. While it is common to decrease the learning rate for smaller batch sizes, other hyperparameters are often held fixed. In this work, we revisit small batch sizes all the way down to batch size one, and we propose a rule for scaling Adam hyperparameters to small batch sizes. In particular, rather than holding the decay rate of the second moment fixed across batch sizes, we propose to hold its half-life fixed in terms of tokens. We find that small batch sizes (1) train stably, (2) are consistently more robust to hyperparameter choices, (3) achieve equal or better per-FLOP performance than larger batch sizes, and (4) notably enable stable language model training with vanilla SGD, even without momentum, despite storing no optimizer state. Building on these results, we provide practical recommendations for selecting a batch size and setting optimizer hyperparameters. We further recommend against gradient accumulation unless training on multiple devices with multiple model replicas. Finally, we show that a small batch size combined with an optimizer with a small state size can provide the performance benefits of full fine-tuning while maintaining a similar memory footprint to LoRA.
@@
Figure 1: Small batch sizes are more robust to optimizer design and hyperparameter values. (a) Loss achieved by a transformer decoder-only language model with 30M active parameters trained on 600M tokens of FineWeb-Edu data using SGD, Adafactor, Adam, and Muon. At small batch sizes, all optimizers achieve similar loss; at large batch sizes, the gap between optimizers grows. (b) Loss achieved by GPT-2 (124M) trained on 2.5B tokens of FineWeb using AdamW when tuning the learning rate and $\beta_{1}$ hyperparameters for batch sizes 1 and 512. While both batch sizes achieve a similar lowest loss, the smaller batch size is much more robust to the hyperparameter values, reducing the need for hyperparameter tuning. For batch size 512, we use the default hyperparameters from Brown et al. (2020) and the nanoGPT public repository (Karpathy, 2023). For batch size 1, we turn off weight decay and rescale $\beta_{2}$ to preserve the token half-life, resulting in $\beta_{2}=0.9999$.
@@ 1 Introduction
Large batch sizes are widely believed to improve the stability of language model training (Gazagnadou et al., 2019; Nitanda et al., 2021; Kim et al., 2023; Ghosal et al., 2023). As a consequence, sophisticated optimizers that perform well in large-batch training are increasingly standard practice (You et al., 2017; Gupta et al., 2018; You et al., 2019; Liu et al., 2023; Jordan et al., 2024; Vyas et al., 2024). In fact, it is common to simulate batch sizes even larger than the maximum batch size that fits into device memory through gradient accumulation (Ghosal et al., 2023; Kim et al., 2023; Zeng et al., 2023; Le Scao et al., 2022).
@@
In small batch size pretraining experiments, one may observe loss spikes and heavy instability Xu et al. (2020); Xiao (2024); Labonne (2024); Weights & Biases (2025). While it is common to decrease the learning rate for smaller batch sizes, other hyperparameters, such as the decay rates for the first and second moments in Adam ($\beta_{1}$ and $\beta_{2}$), are often held fixed across batch sizes. We show that if instead of holding $\beta_{2}$ fixed, we hold the half-life of $\beta_{2}$ fixed (measured in number of tokens), stable training is possible all the way down to batch size one, as illustrated in Figure˜1(a).
@@
When scaling hyperparameters in this way, we find that small batch sizes can confer computational advantages and greater robustness. In the small batch regime, we observe that the speed of convergence is less sensitive to optimizer hyperparameters like the learning rate, and momentum becomes less necessary, as we find in Figure˜1(b). In contrast, large batch sizes require large learning rates to compensate for taking fewer steps at a given compute budget. Large step sizes in turn require that the optimizer make predictions about the loss surface far away from the current iterate. We find that predicting such far-away parameter updates requires a sophisticated and carefully tuned optimizer. For very small batch sizes, we find that even vanilla stochastic gradient descent (SGD), with no momentum or weight decay, becomes competitive for training language models, in contrast to findings from recent work (Zhao et al., 2025). The ability to perform stable training without momentum can confer significant memory advantages, since we do not need to store the corresponding optimizer state, which can have a significant memory footprint. Based on our observations, we also revisit Adafactor, a memory-efficient variant of Adam, and show that it can be a compelling alternative to Adam in the small batch regime, enabling training of larger models in a memory-constrained setting.
@@
These findings could have a significant bearing on practice. Rather than perform gradient accumulation or use the largest batch size that fits into device memory, it could be preferable to use smaller batch sizes that enable simpler procedures, computational and memory advantages, and require less tuning. Our findings suggest a best practice of using the smallest batch size that maximizes model throughput. We find that these prescriptions hold in both LLM pretraining and fine-tuning.
@@
It is common in convex optimization to use higher batch sizes as the loss converges to a minimum (Gazagnadou et al., 2019; Sievert and Shah, 2019). Intuitively, the higher the gradient to gradient noise magnitude ratio, the smaller the batch size we should take. For example, if the gradient noise magnitude is close to zero, then we obtain roughly the same gradient estimate with a small batch size as a large batch size but pay far fewer floating-point operations (FLOPs), so small batch sizes are efficient. Smaller batch sizes allow us to take more steps for a fixed FLOP budget. As we converge to a minimum, that same ratio instead often tends to zero since the gradient norm vanishes while the gradient noise may not, in which case we may want a larger batch size. We hypothesize that language model training lives in the far-from-convergence regime where small batch sizes are efficient. Following compute-optimal scaling laws, we would not train language models into the convergence regime since we could achieve better performance by instead training a larger model on less data (Kaplan et al., 2020; Hoffmann et al., 2022).
@@
The paper is structured as follows: in Sections˜2 and 3, we cover the background and related work for language model training. In Section˜4, we demonstrate that small batch sizes not only perform on par with larger batch sizes for several optimizers that we carefully tune, but they also exhibit more robustness to optimizer and hyperparameter choices. We also show that the gap between vanilla SGD and more sophisticated optimizers shrinks in the small batch regime. Moreover, we derive scaling heuristics for Adam’s hyperparameters and demonstrate that they yield improved performance when transferred to new, larger-scale settings. Finally, we apply these results to memory-efficient fine-tuning using a small batch size and Adafactor, which achieves the best trade-off between memory footprint and performance. Motivated by our findings, we provide practical prescriptions for training language models in Section˜5. Namely, we recommend using the smallest batch size that maximizes model throughput and advise against gradient accumulation for most practitioners.
@@ 2 Background: Optimization for Language Models
Stochastic gradient descent (SGD) (Robbins and Monro, 1951; Bottou, 2010). SGD updates model parameters by computing gradients on mini-batches of data as follows: $\theta_{t+1}=\theta_{t}-\eta g_{t}$, where $\theta_{t}$ are the model parameters at step $t$, $\eta$ is the learning rate, and $g_{t}=\nabla_{\theta}\mathcal{L}(\theta_{t})$ is the gradient of the loss function with respect to $\theta_{t}$ averaged over $B$ samples. We refer to $B$ as the batch size and consider the extreme case of $B=1$ in our experiments.
We show that although SGD is extremely simple, it can perform competitively when correctly tuned at small batch sizes. Momentum is often added to accumulate a moving average of past gradients, smoothing updates, and accelerating convergence. However, we focus on vanilla SGD in our experiments to make the point that, with a small batch size, even momentum is unnecessary.
Adam (Kingma, 2014). Adam is an adaptive optimizer that maintains an exponential moving average (EMA) of the gradient and squared gradient, referred to as the first and second moments, and denoted as $m_{t}$ and $v_{t}$. The timescales of these moving averages are controlled by decay rates $\beta_{1}$ and $\beta_{2}$:
@@
Higher values of $\beta_{1}$ and $\beta_{2}$ place more weight on past gradients, meaning that the moments are averaged over longer timescales and evolve slowly. When training language models, values like $\beta_{1}=0.9$ and $\beta_{2}\in[0.95,0.98]$ are often used in practice (Brown et al., 2020; Touvron et al., 2023). While most researchers and practitioners do not scale $\beta_{1}$ and $\beta_{2}$ with the batch size, we show in our work that scaling $\beta_{2}$ in particular is crucial for achieving good performance with Adam at a small batch size.
@@
Moment half-life. We found it helpful across many of our experiments to measure the typical timescales (measured in number of tokens) that the first and second moments in Adam are averaged over, instead of directly working with $\beta_{1}$ and $\beta_{2}$.
At each Adam update step [eq.˜1], the contribution of previous gradients to the first moment gets reduced by a factor of $\beta_{1}$ and the contribution to the second moment gets reduced by a factor of $\beta_{2}$. Hence, by definition, there exists a certain number of steps $n$, after which the contribution of any mini-batch gradient is halved: $\beta^{n}=\frac{1}{2}$. Since every update step corresponds to observing $(B\cdot T)$ tokens, where $B$ is the batch size and $T$ is the sequence length, we can convert the number of optimizer steps into a number of observed tokens. As such, we can express the number of tokens it takes to decay a mini-batch gradient by a factor of $\frac{1}{2}$ as the half-life of the decay rate, denoted by $t_{\nicefrac{{1}}{{2}}}$, where $\beta^{\frac{t_{\nicefrac{{1}}{{2}}}}{BT}}=\frac{1}{2}$. The half-life provides a measure of the “typical” timescale that gradients are averaged over. For example, if we use a half-life of $t_{\nicefrac{{1}}{{2}}}=10\textrm{M}$ tokens, this means that the second moment is averaged over roughly 10M tokens. Throughout this paper, we will refer to the half-life of the first moment as $t_{1}$ and the half-life of the second moment as $t_{2}$, analogously to the more conventional parameterization $\beta_{1}$ and $\beta_{2}$.
Adafactor (Shazeer and Stern, 2018). The second moment estimate $v_{t}$ used in Adam requires storing a number of floating point values equal to the parameter count of the model itself. Instead, Adafactor only stores the per-row and per-column sums of the moving average for the neural network weight matrices and estimates the per-parameter second moment using these sums. Adafactor also does not store a first moment estimate at all. Therefore, for a weight matrix with shape $d_{1}\times d_{2}$, the memory requirements for the moving averages are reduced from $2\times d_{1}\times d_{2}$ for Adam to $d_{1}+d_{2}$ for Adafactor. We demonstrate in Section˜4.5 that thanks to this sublinear memory cost, Adafactor achieves the best performance–memory trade-off.
Muon (Jordan et al., 2024). Another optimizer that was recently demonstrated to be competitive for language model training is MomentUm Orthogonalized by Newton-Schulz (Muon). Muon is specifically designed for $\geq 2$-dimensional parameter tensors, such as weight matrices in linear layers, where the Newton-Schulz iterative method is applied to the first moment estimate $m_{t}$ before using it to update the parameters. Muon is often used in conjunction with other optimizers like Adam. In our experiments using Muon, we only apply it to hidden layers and use Adam for input embedding and final layers.
@@ 3 Related Work
Batch size has long been studied in both theory and practice: its impact on the optimization path (Jastrzębski et al., 2018; Jastrzebski et al., 2020; Lee et al., 2022; Atanasov et al., 2024), its interaction with different optimizers (Wilson and Martinez, 2003; Chen et al., 2016; Shallue et al., 2019; Zhang et al., 2019), and its effect on the optimal learning rates and momentum hyperparameters (Smith and Le, 2017; Granziol et al., 2022; Wang et al., 2023; Fu et al., 2023; Shallue et al., 2019; Kidambi et al., 2018; Zhang et al., 2024a; Porian et al., 2024; Zhang et al., 2024a; Chowdhery et al., 2023; Zhang et al., 2022). Prior work has also explored related themes, such as the flatness argument favoring SGD (LeCun et al., 2002; Keskar et al., 2016; Goyal et al., 2017; Masters and Luschi, 2018), SGD’s implicit biases (Arora et al., 2019; Barrett and Dherin, 2020; HaoChen et al., 2021; Smith et al., 2021), and the comparison between stochastic and full-batch optimization (Ge et al., 2015; Du et al., 2017; Lewkowycz et al., 2020; Geiping et al., 2021). However, modern language model training is neither convex nor characterized by multiple epoch training like vision settings. In fact, we often train for a limited token budget and stop far before convergence, since otherwise scaling laws suggest that we could train a much better language model for the same compute budget by training a larger model on fewer tokens (Hoffmann et al., 2022). Hence, prior intuitions, such as the need for a larger batch size in the convergence phase, may not apply. For this reason, in the rest of this section, we cover works that focus on modern language model training specifically.
@@
Zhang et al. (2024a) and Shallue et al. (2019) examine the critical batch size, which represents the threshold beyond which greater data parallelism would lead to diminishing returns, and tune all optimizer hyperparameters. They argue that any batch size below the critical batch size should perform equally well when training for the same number of tokens. Differently from these works, we empirically demonstrate that small batch sizes, including the extreme value of batch size one, perform on par with or even better than larger batch sizes, and we also show that small batches are more robust to optimizer and hyperparameter choices and enable SGD to perform competitively. We validate these findings on language models with modern design choices and up to 1.3 billion parameters.
@@
On the other hand, Vyas et al. (2023) argue that gradient noise from small batch sizes does not yield similar performance advantages in the single epoch setting as it does in multiple-epoch training. Experimental results from Filatov et al. (2024) also suggest that small batches do not work as well as large batches for language models trained on C4 (Dodge et al., 2021), but notably, the authors do not tune the decay rates $\beta_{1}$ and $\beta_{2}$ of the Adam optimizer. In the same vein, Xiao (2024) claim that both excessively small and large batch sizes underperform for language model training with Adam, leading to a U-shaped curve of the loss as a function of the batch size. We reproduce the exact experiments performed by Xiao (2024) and show that their conclusions are an artifact of incomplete hyperparameter tuning; we find that simply using a slower decay rate of the second moment estimate $\beta_{2}$ makes small batch sizes competitive.
@@
Zhao et al. (2025) train language models on C4 using different diagonal preconditioning optimizers, namely Adam, Lion, Adafactor with momentum, SGD with momentum, and signed SGD with momentum. They demonstrate that all these optimizers, aside from SGD which lags behind, achieve comparable optimal performance and robustness to hyperparameters. We challenge this conclusion in our work and argue that in the small batch size regime, SGD can perform on par with Adam.
Shallue et al. (2019), Kidambi et al. (2018) and Zhang et al. (2024a) find that momentum is not necessary for SGD when the batch size is small, which aligns with our findings. However, Kidambi et al. (2018) focus on small-scale vision experiments, Shallue et al. (2019) and Kidambi et al. (2018) do not compare the performance of SGD vs. Adam, and Zhang et al. (2024a) claim that SGD performs significantly worse than Adam. We hypothesize this observation could be a result of Zhang et al. (2024a) using 64 as the smallest batch size – we test batch sizes all the way down to 1. Zhang et al. (2019) highlight that the benefits of more sophisticated optimizers diminish as batch size decreases, but their theoretical results assume a convex quadratic objective and their empirical results only cover a single language modeling experiment that is limited to a two-layer transformer.
Porian et al. (2024) and Zhang et al. (2024a) find that increasing the value of $\beta_{2}$ is important for small-batch training with Adam. However, their approach is based on discrete hyperparameter sweeps across $\beta_{2}$. In contrast, we introduce and validate a principled scaling rule that holds the second-moment half-life constant in terms of tokens. Busbridge et al. (2023) propose a scaling rule for the decay coefficient of exponential moving averages similar to ours, but applied in the context of model weight averaging. Zhang et al. (2022) show that there exists a threshold of $\beta_{2}^{*}$ such that when $\beta_{2}$ > $\beta_{2}^{*}$, Adam converges (this threshold decreases with batch size, i.e. the threshold is higher for smaller batch sizes). Chowdhery et al. (2023) note that increasing $\beta_{2}$ during training leads to better performance for rare token embeddings. While this is an important insight, we provide a simpler batch-size-aware heuristic that applies even when using a constant $\beta_{2}$ throughout training.